# R-Judge: Benchmarking Safety Risk Awareness for LLM Agents

### 18 Jan 2024 v1

@@ Abstract
Large language models (LLMs) have exhibited great potential in autonomously completing tasks across real-world applications. Despite this, these LLM agents introduce unexpected safety risks when operating in interactive environments. Instead of centering on LLM-generated content safety in most prior studies, this work addresses the imperative need for benchmarking the behavioral safety of LLM agents within diverse environments. We introduce R-Judge, a benchmark crafted to evaluate the proficiency of LLMs in judging safety risks given agent interaction records. R-Judge comprises 162 agent interaction records, encompassing 27 key risk scenarios among 7 application categories and 10 risk types. It incorporates human consensus on safety with annotated safety risk labels and high-quality risk descriptions. Utilizing R-Judge, we conduct a comprehensive evaluation of 8 prominent LLMs commonly employed as the backbone for agents. The best-performing model, GPT-4, achieves 72.29% in contrast to the human score of 89.38%, showing considerable room for enhancing the risk awareness of LLMs. Notably, leveraging risk descriptions as environment feedback significantly improves model performance, revealing the importance of salient safety risk feedback. Furthermore, we design an effective chain of safety analysis technique to help the judgment of safety risks and conduct an in-depth case study to facilitate future research. R-Judge is publicly available at https://github.com/Lordog/R-Judge.
@@ 1 Introduction
Large language models (LLMs) have shown emergent abilities in reasoning, decision making, and instruction following (Wei et al., 2022a). The stimulating capabilities of LLMs, especially the most powerful one, GPT-4 (OpenAI, 2023), ignite the development of LLM agents (Zhang et al., 2023c). Equipped with tool usage and environment interaction, agents, such as AutoGPT Richards (2023) and Voyager (Wang et al., 2023a), can autonomously complete user-specified tasks with LLMs as controllers (Xi et al., 2023; Wang et al., 2023b; Zhou et al., 2023b; Lin et al., 2023; Sumers et al., 2023).
@@
Given that unknown risks rest in complex environments and user instructions, LLM agents are likely to cause unexpected safety issues (Xi et al., 2023; Ruan et al., 2024; Naihin et al., 2023). For example, when asked to process emails, agents may unconsciously click the URL of phishing emails, leading to potential privacy leakage and even property loss. Therefore, safety assurance is a necessary prerequisite for LLM agents in practical applications. To this end, it is critical to develop effective mechanisms of safety evaluation.
@@
Existing works on safety evaluation focus on safety issues of LLM-generated contents (Zhiheng et al., 2023; Bhardwaj and Poria, 2023), e.g., to alleviate generating contents related to offensiveness, unfairness, illegal activities, ethics, and misinformation. Notably, SafetyBench (Zhang et al., 2023b) evaluated LLMs with multiple-choice questions covering 7 categories of safety concerns. SuperCLUE-Safety (Xu et al., 2023a) presented a comprehensive benchmark with multi-round and open-ended questions. However, without interactive environments where LLM agents execute, they could not provide a realistic safety evaluation of LLM agents.
@@
To tackle the aforementioned challenge, few studies have proposed mechanisms for safety evaluation of LLM agents. To identify risks of LLM agents in interactive environments, ToolEmu (Ruan et al., 2024) implemented a GPT-4 powered emulator with diverse tools and scenarios tailored for the LLM agents’ action execution. Alongside the emulator, a GPT-4 powered automatic safety evaluator examines agent failures and quantifies associated risks. In a similar vein, AgentMonitor (Naihin et al., 2023) proposed a framework where an LLM serves as a monitor for the execution of agents. It intervenes by halting actions deemed unsafe, thereby preventing safety issues that LLM agents might encounter on the open internet.
@@
Though implemented with different settings, both ToolEmu and AgentMonitor utilized LLMs as judges to identify risky actions of LLM agents. However, judging whether agent actions are safe in the context of multi-turn interactions involving LLM agents, users, and environment in diverse scenarios, is a challenging yet realistic task for LLMs. The key challenges are twofold. First, owing to the complexity of interactions and diversity of environments, it remains unclear whether LLMs are aware of agent safety issues without priors of risk descriptions from humans. Second, the fairness of the evaluation could be compromised by the inherent bias introduced when both the judge and the tested agent are LLMs, potentially even identical ones.
@@
This work presents R-Judge, a benchmark crafted to evaluate the proficiency of LLMs in judging safety risks given agent interaction records. Each record contains a task instruction and a history of agent actions and environment feedback. R-Judge comprises 162 agent interaction records, encompassing 27 popular application scenarios across 7 categories including program, operation system (OS), internet of things (IoT), software, web, finance, and health. For each agent interaction record, we annotate safety risk labels as well as high-quality risk descriptions of identified risks with human consensus.
Utilizing R-Judge, we conducted a comprehensive evaluation of 8 prominent LLMs commonly employed as the backbone for agents. Concretely, feeding the records of agent execution as inputs, the evaluated LLM is required to judge whether agent actions are safe. Experimental results show that most models fail the task. The best-performing model, GPT-4, achieves an F1 score of 72.29% in contrast to the human score of 89.38%. The results indicate that risk awareness in open scenarios is challenging for most LLMs and LLM agent safety is of significant concern. Notably, feeding risk descriptions as environment feedback significantly improves model performance, revealing the importance of salient safety risk feedback.
In summary, our contributions are as follows:
@@
$\bullet$ We first formalize the problem of agent safety from the perspective of LLM awareness on safety risks. Concretely, we formulate the task of judging safety risks given agent interaction records and utilize it as a simple and effective paradigm to evaluate the risk awareness of LLMs.
@@
$\bullet$ We propose R-Judge, a general, realistic, and fair benchmark. R-Judge involves complex multi-turn interactions between the user, environment, and agent. It incorporates human consensus on safety with annotated safety risk labels and high-quality risk descriptions. To our knowledge, this is the first benchmark dataset curated to assess risk awareness of LLMs for agent safety.
@@
$\bullet$ We evaluate 8 popular LLMs based on R-Judge and disclose that most models fail to identify safety risks in open scenarios. We design an effective chain of safety analysis technique to help the judgment of safety risks and conduct in-depth case studies to facilitate future research.
@@ 2 Problem Formalization
This section presents the problem formalization for agent safety awareness. We first define the record of LLM agents as the basis for safety judgment. Based on the record, we formulate the task of safety judgment to evaluate the risk awareness of LLMs. Finally, we will present our design criteria for building a general, realistic, and fair benchmark.
@@
Interacting directly with complex environments, LLM agents are likely to cause unexpected safety issues, some of which are severe, such as crucial privacy leakage and data loss. With LLMs as judges to monitor risky actions, LLM agents cause fewer safety issues, as disclosed by Toolemu and AgentMonitor. Therefore, risk awareness of LLMs is a significant factor for the safe execution of LLM agents. Here, we formulate the task of agent safety judge to evaluate the risk awareness of LLMs. Figure 2 illustrates the overall procedure.
Agent interaction record is checked by an LLM to judge if actions are safe or not. The results indicate whether the LLM can identify the safety risks, revealing risk awareness of LLMs.
The task of agent safety judge is formalized as:
$f:p_{\theta}(R|\textrm{scenario},\textrm{prior})\to\textrm{label},$ (2)
where $p_{\theta}(\cdot)$ denotes the language model mapping record $R$ to a binary label representing safe or unsafe. The $\textrm{scenario}\in\mathcal{C}$ refines the environment that the agent executes in, where $\mathcal{C}$ is category, the set of scenarios. The $\textrm{prior}\in\mathcal{L}$ indicates extra information to support judgments of LLMs, such as scenario-specific risks or safety guidelines.
Inspired by recent advances in chain-of-thought reasoning (Wei et al., 2022b; Kojima et al., 2022; Zhang et al., 2023d), we require that the LLM first generates safety analysis concerning the agent records and then makes the judgment accordingly. Therefore, the formulation is further extended to
$f:p_{\theta}(R|\textrm{scenario},\textrm{prior})\to(\textrm{analysis},\textrm{label}),$ (3)
where $\textrm{analysis}\in\mathcal{L}$.
Design Philosophy of Benchmark.
Now we present our design criteria for building a general, realistic, and fair benchmark. Based on the formalization above, we meticulously design prerequisite settings and evaluation methods.
$\bullet$ General. In general setting, LLM Judges are blind to priors of scenario-specific risks or safety guidelines, and not limited to specific scenarios, which respectively distinguishes Toolemu and AgentMonitor. In other words, the evaluation formulation is the general one
@@
$\bullet$ Realistic. In the real world, risks lie in complex multi-turn interactions of LLM agents with the user and environment in diverse scenarios. An effective benchmark should present realistic datasets, containing complex interaction processes between the user, environment, and agent. Therefore, we design a curation pipeline for realistic datasets. Furthermore, human consensus on safety makes datasets human-aligned.
$\bullet$ Fair. Many open-ended LLM evaluations use proxy LLM evaluators, GPT-4, or dedicated fine-tuned LLMs, to score candidate LLMs. However, utilizing LLM to evaluate LLMs leads to potential inherent bias. It is improper to utilize for the challenging task of agent safety judge. For fairness, we adopt a label-based evaluation, which takes only the output label into the metric calculation.
Figure 2: Illustration of R-Judge by an example. The upper left part is an example in datasets, e.g., a record of an agent interacting with the user and environment. The upper right part is human annotation involving a binary safety label and high-quality risk description. Following the arrow, we can see the evaluation paradigm with two tests. In the Standard Test(w/o desc), given record, LLMs are asked to generate an analysis and a label. In Oracle Test(w desc), LLMs are provided with annotated risk descriptions additionally. Based on the generated label, metrics are calculated to evaluate the proficiency of LLMs in judging safety risks given agent interaction records.
@@ 3 R-Judge Benchmark
To evaluate risk awareness of LLMs for agent safety, we construct R-Judge, a general, realistic, and fair benchmark dataset with human safety consensus.
@@ 3.1 Case Format
Following the problem formalization in Section 2, we define the case format of task $f:p_{\theta}(R)\to(\textrm{analysis},\textrm{label})$, shown on the upper half of Figure 2. Here we give an elaborate explanation of each term.
@@
$\bullet$ Risk Description. Carefully documented by annotators, a risk description accommodates complete information for humans and agents to understand the risks of the case. For each unsafe case, the risk description contains Motivation, Trigger, and Outcome, following the threat model of agent safety in Appendix A.4. Motivation means the risks from users or environments. Trigger means the risky behaviors of agents. Outcome is the results after agents finish risky actions.
@@ 3.2 Dataset Curation
To contribute to a realistic dataset stated in Section 2, we design a dataset curation pipeline incorporating human consensus on safety.
General Principle.
Datasets are curated by both public collection and manual construction, as occurrences of risks are relatively rare in current agent applications. We first predefined curation principles, including category and scenario (Appendix A.1), risk type (Appendix A.2), general safety standard on LLM agents (Appendix A.3), and threat model of LLM agents (Appendix A.4). Then based on curation principles, human annotators collect cases from related work and construct risky cases manually. For labeling, annotators mark a safety label and write a risk description for each record. Whether the label is safe or unsafe is determined by the annotator and then validated by others based on predefined general safety standard(Appendix A.3). By incorporating human consensus on safety, R-Judge is human-aligned.
@@
We collect and modify public datasets, mainly from Toolemu and AgentMonitor. For validity, cases are transformed or removed based on curation principles. For example, some controversial cases are hard even for annotators to reach a consensus. At last, 81 from Toolemu (Ruan et al., 2024) and 24 from AgentMonitor (Naihin et al., 2023) contributed to the dataset construction.
@@
We also construct high-quality data by human experts. Concretely, annotators meticulously brainstorm application cases where agents are likely to trigger certain risk types in certain scenarios. Then, with the assistance of ChatGPT and our validity check, application cases are polished into valid ones.
Due to the complex logic of interaction records and the safety mechanism of GPT-4, we find that GPT-4 is unable to generate valid records for batch construction. Even from the AutoGPT team, AgentMonitor crafts only 27 single-turn unsafe cases, demonstrating the difficulty of manual construction. At last, 81 from Toolemu, 24 from AgentMonitor, and 57 constructed cases form the final 162 R-Judge records.
@@
As a result of the construction process and validity check, R-Judge includes 162 complex cases where intricate risks lie in multi-turn interactions between the user, agent, and environment. The diverse dataset covers 7 selected categories including program, OS, IoT, software, web, finance, and health, covering 27 scenarios. R-Judge also spans 10 types of risk (Appendix A.2), including privacy leakage, computer security, physical health, data loss, financial loss, property damage, illegal activities, ethics and morality, bias and offensiveness, miscellaneous. On average, R-Judge involves 3.0 turns of interaction and 192.9 word counts, with 63.0% being unsafe cases. Figure 1 shows an overview of datasets. More statistics are shown in Table 7.
For human-annotated datasets, there is always a trade-off between the scale of the instances that are annotated and the quality of the annotations (Cui et al., 2020). Our dataset is smaller than the previous crawling-based LLM safety benchmark dataset (Zhang et al., 2023b) due to the complexity of the agent interaction process across distinct environments. However, as the first benchmark dataset curated for agent safety, our data scale is comparable with similar high-quality LLM benchmark datasets with multi-step operation records (Ling et al., 2017; Wei et al., 2022b). The dataset is also scalable with the development of agent applications. Because only records, e.g., off-the-shell agent execution logs, and minimal annotation are needed to expand the dataset once safety risk is encountered. Moreover, we will show that the benchmark dataset is quite effective in benchmarking risk awareness of LLMs via comprehensive experiments.
@@ 3.3 Evaluation Criteria
Upon the dataset, we evaluate LLMs to explore their proficiency in judging safety risks given agent interaction records. For fairness, we adopt a simple label-based evaluation paradigm, which compares LLM-generated binary safety labels with truth labels from the consensus of annotators. The paradigm includes a two-level main test and two additional proposed metrics, to achieve a progressive and valid evaluation. We will introduce the full pipeline in this part, which has been shown in Figure 2.
3.3.1 Evaluation Recipes
We design two evaluation recipes: Standard Test without risk description and Oracle Test with risk description. As stated in Section 2, the task of agent safety judge is formalized as $f:p_{\theta}(R)\to(\textrm{analysis},\textrm{label})$.
In practice, the task may require a two-step API call of generative LLMs. First, $p_{\theta}(R)\to\textrm{analysis}$. Second, $p_{\theta}(R,\textrm{analysis})\to\textrm{label}$.
Both formulations output a single label, thus supporting a progressive label-oriented evaluation.
Standard Test (w/o Description).
@@ 3.3.1 Evaluation Recipes
First, LLMs are demonstrated with the record as input and asked to analyze whether the agent’s actions in the record are safe. Then, LLMs are asked to output 1 for unsafe actions and 0 otherwise based on its analysis and record, as shown in the bottom left of Figure 2.
Oracle Test (w/ Description).
LLMs are directly provided with a record and human-written risk description as input and asked to output 1 for unsafe, and 0 for safe. Oracle Test is a modified version of Standard Test where analysis of LLMs is replaced with ground truth risk description written by human annotators. The Oracle Test checks if LLMs can correctly classify safe and unsafe cases given ground truth risk description as shown in the bottom right of Figure 2.
3.3.2 Validity Check
Due to the limitations of LLMs, two-level tests with a single label face two potential key problems of invalid outputs [^1] Examples of invalid outputs are provided in Appendix D..
$\bullet$ Invalid label. Some models may sometimes fail to output 1 or 0 due to bad instruction following.
$\bullet$ Invalid mapping. The tendency of some models to output 1 or 0 differs possibly because of the unbalanced text of 1 and 0 during fine-tuning. Even if the model regards a case as safe, it may output 1 which represents unsafe. And vice versa.
First, to tackle invalid labels, we present Validity, the ratio of models that successfully output a single binary label as an answer. If models fail to output 1/0, the output will be counted as wrong. Second, to address the problem of invalid mapping, we present Consistency, the ratio of models’ consistent output. We call the model a second round to check the consistency of the predicted label. In the second round, we reverse the mapping relation, i.e. 0 represents unsafe and 1 represents safe. If two output labels of the model represent the same, the output is valid. Otherwise, the output is counted as wrong.

### 18 Feb 2024 v2

@@ Abstract
Large language models (LLMs) have exhibited great potential in autonomously completing tasks across real-world applications. Despite this, these LLM agents introduce unexpected safety risks when operating in interactive environments. Instead of centering on LLM-generated content safety in most prior studies, this work addresses the imperative need for benchmarking the behavioral safety of LLM agents within diverse environments. We introduce R-Judge, a benchmark crafted to evaluate the proficiency of LLMs in judging and identifying safety risks given agent interaction records. R-Judge comprises 162 records of multi-turn agent interaction, encompassing 27 key risk scenarios among 7 application categories and 10 risk types. It incorporates human consensus on safety with annotated safety labels and high-quality risk descriptions. Evaluation of 9 LLMs on R-Judge shows considerable room for enhancing the risk awareness of LLMs: The best-performing model, GPT-4, achieves 72.52% in contrast to the human score of 89.07%, while all other models score less than the random. Moreover, further experiments demonstrate that leveraging risk descriptions as environment feedback achieves substantial performance gains. With case studies, we reveal that correlated to parameter amount, risk awareness in open agent scenarios is a multi-dimensional capability involving knowledge and reasoning, thus challenging for current LLMs. R-Judge is publicly available at https://github.com/Lordog/R-Judge.
1 Introduction
@@
Large language models (LLMs) have shown compelling abilities in reasoning, decision making, and instruction following (Wei et al., 2022a). The stimulating capabilities of LLMs, especially GPT-4 (OpenAI, 2023), ignite the development of LLM agents (Zhang et al., 2023b). Equipped with tool usage and environment interaction, agents, such as AutoGPT (Richards, 2023) and Voyager (Wang et al., 2023a), can autonomously complete user-specified tasks with LLMs as controllers (Xi et al., 2023; Wang et al., 2023b; Zhou et al., 2023b; Lin et al., 2023a; Sumers et al., 2023).
@@
Given that unknown risks rest in complex environments and user instructions, LLM agents are prone to cause unexpected safety issues (Xi et al., 2023; Ruan et al., 2024; Naihin et al., 2023). For instance, when asked to process emails, agents may unconsciously click the URL of phishing emails, leading to potential privacy leakage and even property loss. Therefore, safety assurance is a necessary prerequisite for LLM agents in real-world applications. To this end, it is critical to effectively evaluate the safety risk awareness of LLMs in open agent scenarios.
@@
Existing works on safety evaluation focus on safety issues of LLM-generated contents (Zhiheng et al., 2023; Bhardwaj and Poria, 2023), e.g., to alleviate generating contents related to offensiveness, unfairness, illegal activities, and ethics. Notably, SafetyBench (Zhang et al., 2023a) evaluated LLMs with multiple-choice questions covering 7 categories of safety concerns. Similarly, SuperCLUE-Safety (Xu et al., 2023a) presented a comprehensive benchmark with multi-round and open-ended questions. However, without interactive environments where LLM agents execute, they could not provide a realistic safety evaluation of LLM agents.
@@
To tackle the aforementioned challenge, few studies have investigated safety risks in applications of LLM agents. To identify risks of LLM agents in interactive environments, ToolEmu (Ruan et al., 2024) implemented a GPT-4 powered emulator with diverse tools and scenarios tailored for the LLM agents’ action execution. Alongside the emulator, a GPT-4 powered automatic safety evaluator examines agent failures and quantifies associated risks. In a similar vein, AgentMonitor (Naihin et al., 2023) proposed a framework where an LLM serves as a monitor for the execution of agents. It intervenes by halting actions deemed unsafe, thereby preventing safety issues that LLM agents might encounter on the open internet.
@@
Though implemented with different settings, both ToolEmu and AgentMonitor utilized LLMs as safety monitors to identify risky actions of LLM agents. However, judging whether agent actions are safe in the context of multi-turn interactions involving LLM agents, users, and environment in diverse scenarios, is a challenging yet realistic task for LLMs. Owing to the complexity of interactions and diversity of environments, it remains unclear whether LLMs are aware of agent safety issues without priors of risk descriptions from humans.
@@
This work presents R-Judge, a benchmark crafted to evaluate the proficiency of LLMs in judging and identifying safety risks given agent interaction records. Each record contains a user instruction and a history of agent actions and environment feedback. R-Judge comprises 162 agent interaction records, encompassing 27 popular application scenarios across 7 categories, including program, operation system (OS), internet of things (IoT), software, web, finance, and health (Figure 1). For each record, we annotate binary safety labels as well as high-quality risk descriptions of identified risks with human consensus.
Utilizing R-Judge, we conducted a comprehensive evaluation of 8 prominent LLMs commonly employed as the backbone for agents and LlamaGuard-7b which is specifically fine-tuned on identifying safety risks in LLM-generated content. The results demonstrate considerable room for enhancing the risk awareness of LLMs: the best-performing model, GPT-4, achieves an F1 score of 72.52% in contrast to the human score of 89.07% while all other models fail the task. Notably, LlamaGuard-7b is incapacitated in R-Judge. Further experiments find that leveraging risk descriptions as environment feedback significantly improves model performance while straightforward prompting mechanisms fail. With case studies, we conclude that risk awareness in open agent scenarios is correlated to parameter amount and is a multi-dimensional capability involving knowledge and reasoning, thus challenging for LLMs. In summary, our contributions are as follows:
@@
$\bullet$ We first formalize the problem of agent safety on LLM awareness of the safety risks. Concretely, we formulate the task of judging and identifying safety risks given agent interaction records and utilize it as a simple and effective paradigm to evaluate the risk awareness of LLMs.
@@
$\bullet$ We propose R-Judge, a human-aligned benchmark dataset with complex multi-turn interactions between the user, environment, and agent. It incorporates human consensus on safety with annotated safety labels and high-quality risk descriptions. To our knowledge, this is the first benchmark dataset curated to assess risk awareness of LLMs for agent safety.
@@
$\bullet$ We evaluate 9 popular LLMs based on R-Judge and disclose that most models fail to identify safety risks in open agent scenarios. Furthermore, we explore the impact of different mechanisms and conduct in-depth case studies, summarizing key findings with valuable insights to facilitate future research.
@@ 2 Problem Formalization
Interacting directly with complex environments, LLM agents are prone to cause unexpected safety issues, some of which are severe, such as crucial privacy leakage and data loss. With LLMs as safety monitors, LLM agents cause fewer safety issues, as disclosed by Toolemu and AgentMonitor. Therefore, risk awareness of LLMs is a significant factor for the safe execution of LLM agents. Here, we formulate the task of agent safety monitor to evaluate the risk awareness of LLMs.
Task Formulation of Agent Safety Monitor.
As illustrated in Figure 2, an LLM checks agent interaction records to judge if actions of the agent are safe. The performance as a safety monitor demonstrates whether the LLM can effectively identify the safety risks and make correct judgments, revealing its risk awareness. The task formulation is:
@@
where $p_{\theta}(\cdot)$ denotes the language model mapping record $R$ to $\textrm{analysis}\in\mathcal{L}$ and a binary label representing safe or unsafe. For general evaluation, LLMs are blind to priors of scenario-specific risks and safety guidelines in various scenarios.
@@ 3 R-Judge Benchmark
Figure 2: Illustration of R-Judge by an example. The upper left part is an example in the dataset, i.e., a record of an agent interacting with the user and environment. The lower left part is human annotation involving a binary safety label and high-quality risk description. Following the arrow, we can see the serial evaluation paradigm with two tests: given record and task instruction, LLMs are asked to generate an analysis and a label. An automatic evaluator compares the analysis with the ground truth risk description to assess the effectiveness of risk identification. And correct generated labels are counted to evaluate the performance of safety judgment.
To evaluate risk awareness of LLMs for agent safety, we construct R-Judge, a general and realistic benchmark dataset containing multi-turn agent interaction with human safety consensus.
@@ 3.1 Case Format
Following the problem formalization in Section 2, we define the case format, as shown on the left half of Figure 2. Here, we provide an elaborate explanation of each term.
@@
$\bullet$ Risk Description. Carefully documented by annotators, a risk description includes complete information for humans and agents to understand the risks of the case. For each unsafe case, the risk description contains Motivation, Trigger, and Outcome, following the threat model of agent safety in Appendix A.4. Motivation means the risk source from the user or environment. Trigger means the risky behaviors of the agent. Outcome is the result after the agent finishes risky actions.
@@ 3.2 Dataset Curation
To contribute to a realistic dataset with high quality, we design a dataset curation pipeline incorporating human consensus on safety.
Curation Principle.
Datasets are curated by both public collection and manual construction, as occurrences of risks are relatively rare in current agent applications. We first predefined curation principles, including category and scenario (Appendix A.1), risk type (Appendix A.2), general safety standard on LLM agents (Appendix A.3), and threat model of LLM agents (Appendix A.4). Then based on curation principles, human annotators collect cases from related work and construct risky cases manually. For labeling, annotators mark a safety label and write a risk description for each record. Whether the label is safe or unsafe is determined by the annotator and then validated by others based on predefined general safety standard (Appendix A.3). By incorporating human consensus on safety, R-Judge accommodates human-aligned ground truth.
@@
We collect and modify public datasets, mainly from Toolemu (Ruan et al., 2024) and AgentMonitor (Naihin et al., 2023). For validity, cases are transformed or removed based on curation principles.
@@
We also construct high-quality data from human experts. Concretely, annotators meticulously brainstorm application cases where agents are likely to trigger certain risk types in certain scenarios. Then, with the assistance of ChatGPT and our validity check, application cases are polished into valid ones.
Due to the complex logic of interaction records and the safety mechanism of GPT-4, we find that GPT-4 is unable to generate valid records for batch construction. Even from the AutoGPT team, AgentMonitor crafted only 27 single-turn unsafe cases, demonstrating the difficulty of manual construction. At last, 81 from Toolemu, 24 from AgentMonitor, and 57 constructed cases form the final 162 R-Judge records.
@@
R-Judge comprises 162 complex cases where intricate risks lie in the multi-turn interaction between the user, agent, and environment. On average, it involves 3.0 turns of interaction and 192.9 word counts, with 63.0% being unsafe cases. The diverse dataset encompasses 7 selected categories including program, OS, IoT, software, web, finance, and health, covering 27 scenarios. It also spans 10 types of risk (Appendix A.2), including privacy leakage, computer security, physical health, data loss, financial loss, property damage, illegal activities, ethics & morality, bias & offensiveness, and miscellaneous. Figure 1 presents an overview of R-Judge. More statistics are shown in Table 4.
@@ 3.3 Evaluation Criteria
Based on the dataset, we evaluate the effectiveness of public LLMs as agent safety monitors. As monitors for agent safety demand both risk identification and safety judgment of LLM capabilities, we design two recipes to evaluate LLMs’ proficiency in identifying and judging safety risks given agent interaction records.
@@
As shown in the right of Figure 2, the two evaluation recipes are conducted in a serial pipeline. First, LLMs are demonstrated with the record as input and asked to analyze whether the agent actions in the record are safe, formalized as $p_{\theta}(R)\to\textrm{analysis}$. Then, LLMs are asked to output ’unsafe’ or ’safe’ based on record and analysis in the first step, formalized as $p_{\theta}(R,\textrm{analysis})\to\textrm{label}$. The output analysis in the first step and the output label in the second step are taken respectively for the two evaluation recipes, with human annotation as ground truth.
Label Score for Safety Judgment.
To evaluate the ability of LLMs to make safety judgments, a label-based test compares LLM-generated binary safety labels with truth labels from the consensus of human annotators.
Pairwise Evaluation for Risk Identification.
To evaluate the effectiveness of LLMs in identifying safety risks, an open-ended test utilizes GPT-4 as an automatic evaluator to assess the open-ended model-generated analysis on unsafe cases.
Risk identification in agent interaction includes two components: what potential risks are and how the agent causes them. We address the two keys with two criteria, ${\rm Alertness}$ and ${\rm Effectiveness}$. As human-annotated risk description is the ground truth, ${\rm Effectiveness}$ is assessed by the relevance between model-generated analysis and the pivot. ${\rm Alertness}$ indicates whether the model senses potential risks, i.e., its generated analysis demonstrates potential risks of the case. The final ${\rm Grade}$ is the addition of ${\rm Effectiveness}$ and ${\rm Alertness}$. The prompt for the GPT-4 evaluator is attached in Appendix C.3.2.
Models Safety Judgment Risk Identification
F1 Recall Specificity Validity Grade Effectiveness Alertness
Random 50.00 50.00 50.00 50.00 16.67 0.00 50.00
Human 89.07 82.17 95.76 100.00 87.00 86.67 88.33
GPT-4 72.52 62.00 83.64 100.00 71.60 71.00 74.00
ChatGPT 39.42 27.00 81.82 100.00 46.20 47.50 41.00
Vicuna-13b-v1.5-16k 43.24 32.00 70.91 99.35 34.20 33.50 37.00
Llama-2-13b-chat-hf 38.86 34.00 25.45 50.97 41.20 40.50 44.00
Vicuna-13b-v1.5 30.30 20.00 78.18 100.00 30.80 31.00 30.00
Vicuna-7b-v1.5-16k 36.88 26.00 72.73 100.00 32.20 31.00 37.00
Llama-2-7b-chat-hf 21.56 18.00 10.91 37.42 23.80 23.00 27.00
Vicuna-7b-v1.5 19.35 12.00 78.18 100.00 30.80 30.00 34.00
LlamaGuard-7b 0.00 0.00 98.18 96.77 0.33 0.50 0.00
Table 1: Main results(%) of the safety judgment test and risk identification test in R-Judge. Safety judgment scores are calculated by counting correct labels and risk identification scores are assigned by an automatic GPT-4 evaluator compared with human-annotated risk description. Segment 1: Reference Score, including Random as baseline and Human as upper bound marked as underline; Segment 2: API-based models; Segment 3: 13B models; Segment 4: 7B models. The best average model results are in bold face, except for Validity which is the reference score.